OpenStudy (anonymous):

What is the lim of h approaching 0 of (3^2-(3+h)^2)/(h) and how do you solve it?

zepdrix (zepdrix):

Was the problem presented like this? Because your numerator looks backwards.

zepdrix (zepdrix):

\[\Large\rm \lim_{h\to0}\frac{(3+h)^2-3^2}{h}\]This would be the limit definition of the derivative of the function \(\Large\rm f(x)=x^2\) evaluated at x=3.

OpenStudy (anonymous):

\[\frac{ 3^{2}-(3+h)^{2} }{ h }\]

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\]

zepdrix (zepdrix):

Alright fine :O If you say so... So expand out your (3+h)^2. Remember how to do that?\[\Large\rm (3+h)^2=(3+h)(3+h)=?\]

OpenStudy (anonymous):

so that would =9+6h+h^2

zepdrix (zepdrix):

\[\Large\rm \lim_{h\to0}\frac{3^2-(9+6h+h^2)}{h}\]Ok good. Understand why I put it in brackets? It's important that you distribute the subtraction to each of the three terms.

OpenStudy (anonymous):

yes, so i have changed the signs to \[\lim_{h \rightarrow 0}\frac{ 3^{2}+9-6h-h ^{2} }{ h }\] that do i do with the \[3^{2}\]

zepdrix (zepdrix):

Square the 3 silly 0_o

zepdrix (zepdrix):

Woops you didn't distribute the minus to the 9.

OpenStudy (anonymous):

Oh woops, so the 9's would cross out leaving \[\frac{ -6h-h ^{2} }{ h }\] and i understand how to find the rest of the problem. Thank you very much!!

zepdrix (zepdrix):

yay team!